ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 145, 2005

Astashkin, S. V. and F. A. Sukochev Series of independent random variables in rearrangement invariant spaces: An operator approach

125

Avkhadiev, F. G. and K.-J. Wirths The conformal radius as a function and its gradient image

349

Baruch, E. M. and Z. Mao Bessel identities in the Waldspurger correspondence over the real numbers

1

Chin, W. and L. Krop Quantized hyperalgebras of rank 1  193

Dernek, N. and O. A. Veliev On the Riesz basisness of the root functions of the nonself-adjoint Sturm–Liouville operator

113

Deville, R. and P. Hájek On the range of the derivative of Gâteauxsmooth functions on separable Banach spaces

257

Ebrahimian, R., D. Kiani and M. Mahdavi-Hezavehi Supersoluble crossed product criterion for division algebras

325

Foias, C., I. B. Jung, E. Ko and C. Pearcy Operators that admit a moment sequence

83

Gardiner, S. J. and A. Gustafsson Pointwise convergence and radial limits of harmonic functions

243

Givon, D., O. H. Hald and R. Kupferman Existence proof for orthogonal dynamics and the Mori–Zwanzig formalism

221

Goren, A. A measuring argument for finite permutation groups  333

Gustafsson, A. See S. J. Gardiner

Hájek, P. See R. Deville

Hald, O. H. See D. Givon

Hinrichs, A. and C. Richter The Knaster problem: More counterexamples

311

Jung, I. B. See C. Foias

Kiani, D. See R. Ebrahimian

Ko, E. See C. Foias

Krop, L. See W. Chin

Kupferman, R. See D. Givon

Mahdavi-Hezavehi, M. See R. Ebrahimian

Mao, Z. See E. M. Baruch Mináč, J. and J. Swallow Galois embedding problems with cyclic

quotient of order p

93

Pearcy, C. See C. Foias Pestov, V. A corrigendum to the article: Ramsey–Milman pheno-

menon, Urysohn metric spaces, and extremely amenable groups

375

Plagne, A. and W. A. Schmid On large half-factorial sets in elementary p-groups: Maximal cardinality and structural characterization

285

Repovš, D., M. Skopenkov and F. Spaggiari On the Pontryagin– Steenrod–Wu theorem

341

Richter, C. See A. Hinrichs Schmid, W. A. See A. Plagne Singh, A. and M. Thakur Reality properties of conjugacy classes

in G_{2} 157

Skopenkov, M. See D. Repovš

Solanes, G. Integral geometry of equidistants in hyperbolic space  271

Spaggiari, F. See D. Repovš Sukochev, F. A. See S. V. Astashkin Swallow, J. See J. Mináč Thakur, M. See A. Singh Veliev, O. A. See N. Dernek Wirths, K.-J. See F. G. Avkhadiev